# Diffusion Models for Time Series Forecasting: A Survey

### v1

@@ 8 Future Directions
Despite recent advances, diffusion-based TSF continues to face open challenges related to generalization, adaptability, and scalability. These issues suggest several promising avenues for future research, outlined below.
Foundation Models
One important direction involves the development of general-purpose time series foundation models that can operate across diverse application domains. Such models are expected to integrate self-supervised pretraining and retrieval-augmented memory, enabling robust few-shot and zero-shot generalization. By learning unified latent representations that align numerical sequences, textual descriptions, and spatial context, these models significantly reduce dependence on domain-specific labels and manual adaptation.
Adaptive Architectures
Another promising direction is the design of adaptive model architectures that automatically adjust key forecasting parameters, such as the historical context length and the prediction horizon, based on the characteristics of the input. This adaptability allows a single model to generalize across various forecasting settings without retraining. Potential solutions include hierarchical attention mechanisms, context-aware gating, and inference-time scheduling strategies.
Long-term Multivariate Forecasting
Forecasting over extended horizons in high-dimensional multivariate settings remains a major technical challenge. Future models should incorporate explicit mechanisms to mitigate error accumulation. In addition, it is necessary to design efficient architectures that can effectively capture cross-variable dependencies with reduced computational cost. Another promising approach is the incorporation of auxiliary multimodal priors, including textual reports, spatial data, or remote sensing imagery. These complementary information sources enhance forecast stability, interpretability, and robustness in complex real-world domains such as energy management, climate prediction, and financial risk analysis.

### v2

@@ 8 Future Directions
Diffusion-based TSF continues to advance along several complementary axes. The directions below focus on scalable pretraining, interpretable uncertainty, trajectory-level multimodal fusion, process optimization for temporal structure, reduction of inference cost, and broader application coverage, with an emphasis on concrete mechanisms and evaluation protocols that match real deployment settings.
8.1 Foundation diffusion models for universal TSF
A foundation diffusion model for time series pretrains a generative prior on heterogeneous corpora spanning domains, sampling regimes, and horizon specifications. The objective covers forecasting under diverse history lengths, variable sets, and horizons, together with multi view masking, reconstruction across gaps, trajectory infilling at multiple scales, and contrastive alignment across datasets. Regularization promotes invariance to time shift, amplitude scaling, and calendar effects. Representation sharing separates domain invariant dynamics from domain specific signals so that the prior transfers across tasks with minimal additional training. Parameter efficient adaptation supports specialization through lightweight adapters, prompt-based conditioning, and low rank updates that preserve the pretrained prior while aligning to a target dataset. Retrieval augmented conditioning supplies similar histories or prototypical patterns from large repositories. Retrieved sequences act as structured conditions that guide score estimation and reduce brittleness under distribution shift and data scarcity. Practical deployment benefits from continual pretraining with drift detection, from domain routers that activate experts under new regimes, and from governance of data provenance. A single backbone with these capabilities sustains robustness under drift and reduces per-task retraining.
8.2 Explainable uncertainty for risk prediction
Future work on uncertainty targets statistical reliability together with transparent attribution. Noise schedules and covariance structures tie to history dependent states and external drivers so that uncertainty tracks volatility cycles, regime shifts, and detected structural breaks. Predicted quantiles form time-indexed curves that remain ordered. At each step the upper quantiles stay above the lower ones and the curves do not cross. Prediction intervals widen as the confidence level increases across the horizon, which preserves internal consistency. Attribution maps connect interval width to trend change, seasonal phase, exogenous events, and sensor quality, with step wise summaries that trace how risk propagates across the horizon. Evaluation report coverage by horizon, by context, and by event type, rather than only aggregate scores. Domain knowledge enters through soft constraints on the score or regularization of the scheduler, which suppresses trajectories that violate physical or operational limits. The outcome is a forecast that reports values, uncertainty, and reasons, with uncertainty that is interpretable and traceable across time.
8.3 Diffusion centric multimodal fusion
Multimodal fusion advances beyond side features by embedding auxiliary signals directly into the diffusion trajectory. Text, spatial or relational graphs, images, and knowledge sources influence the forward prior, the reverse drift, and the latent covariance so that external context shapes generation at each step. Temporal asynchrony receives explicit treatment through alignment operators that account for lags and variable sampling rates during both training and inference. Streaming settings benefit from memory modules that revise beliefs when a delayed modality arrives and from schedules that simulate realistic delays. Reliability varies across sources, therefore confidence weighting becomes a first class component that modulates both the mean and the dispersion. Confidence signals derive from self-assessment of each modality encoder, from agreement between sources, and from historical calibration of attribution errors. Training curricula include lag perturbation, modality dropout, and conflict resolution so that the fusion policy remains stable under missing or noisy inputs. This trajectory level design delivers deeper fusion and raises robustness when the history window provides weak evidence.
8.4 Process optimization for temporal structure
Forward and reverse dynamics align more closely with properties of time series. Schedulers adapt across scales so that coarse trend and fine seasonal components receive appropriate step sizes. Decomposition-aware designs route trend, seasonality, and residual through coupled paths that exchange information during denoising. Frequency-aware latents represent periodic structure in the spectrum while time domain latents retain local transitions, with cross domain coupling that preserves phase information. Time warping maps irregular sampling to a stationary frame and returns forecasts to the original clock, which stabilizes training when intervals vary. Causality respecting trajectories restrict information flow to past and present conditions and enforce lag structure across variables. These refinements produce trajectories that reflect trend, seasonality, non stationarity, and dependence.
8.5 Lowering inference cost
Efficient forecasting with diffusion requires fewer steps and better systems integration. Learned solvers and consistency objectives distill long schedules into short sampling paths while preserving calibration and cross variable dependence. Horizon-aware step allocation places more steps where dynamics change rapidly and fewer steps where signals remain stable. State caching and shared computation across output channels reduce redundant score evaluations. Operator fusion, mixed precision, quantization aware training, and memory efficient attention lower latency and memory footprint. Early exit rules stop sampling when mean and dispersion stabilize under predefined tolerances with guarantees that respect coverage targets. These choices expand the feasible range of horizons and dimensionalities under real time budgets and improve the reproducibility of reported costs.
8.6 Application expansion
Diffusion for time series extends to domains with irregular sampling, exogenous shocks, and rich context. Finance requires event aware volatility tracking, strict auditability of uncertainty, and stress tests that reflect liquidity and market microstructure. Healthcare requires alignment between clinical notes and physiological signals, robust handling of missingness, and safeguards that protect patient privacy. Industrial monitoring benefits from graph informed priors over sensor networks, early warnings for regime changes, and attribution that localizes faults. Earth observation and renewable energy introduce large scale spatiotemporal structure with external drivers from meteorology and policy. Human mobility and operations management require integration of control signals and resource constraints. Progress in these areas depends on datasets and protocols that reflect streaming input, delayed modalities, and deployment constraints, together with metrics that assess accuracy, calibration, attribution quality, and decision utility.